# Theorem Prover as a Judge for Synthetic Data Generation

### v1

@@ Abstract
The demand for synthetic data in mathematical reasoning has increased due to its potential to enhance the mathematical capabilities of large language models (LLMs). However, ensuring the validity of intermediate reasoning steps remains a significant challenge, affecting data quality. While formal verification via theorem provers effectively validates LLM reasoning, the autoformalisation of mathematical proofs remains error-prone. In response, we introduce iterative autoformalisation, an approach that iteratively refines theorem prover formalisation to mitigate errors, thereby increasing the execution rate on the Lean prover from 60% to 87%. Building upon that, we introduce Theorem Prover as a Judge (TP-as-a-Judge), a method that employs theorem prover formalisation to rigorously assess LLM intermediate reasoning, effectively integrating autoformalisation with synthetic data generation. Finally, we present Reinforcement Learning from Theorem Prover Feedback (RLTPF), a framework that replaces human annotation with theorem prover feedback in Reinforcement Learning from Human Feedback (RLHF). Across multiple LLMs, applying TP-as-a-Judge and RLTPF improves benchmarks with only 3,508 samples, achieving 5.56% accuracy gain on Mistral-7B for MultiArith, 6.00% on Llama-2-7B for SVAMP, and 3.55% on Llama-3.1-8B for AQUA.
@@ 1 Introduction
Recent studies have shown that large language models (LLMs) are increasingly capable of tackling mathematical problems Dubey et al. (2024); Jaech et al. (2024); Yang et al. (2024a); Guo et al. (2025). Techniques such as Chain of Thought (CoT) aim to enhance the reasoning process by breaking complex problems into intermediate steps Wei et al. (2022); Kojima et al. (2022). Although these methods have achieved success on benchmarks like GSM8K Cobbe et al. (2021), they are prone to critical flaws. Specifically, LLM-generated CoT reasoning steps may often include errors, leading to incorrect solutions even when the final answer appears plausible Leang et al. (2024).
@@
Parallel to these strategies, synthetic data generation emerged as a promising approach to enhance mathematical reasoning Xin et al. (2024); Lu et al. (2024b). A commonly used verification method is LLM-as-a-Judge Zheng et al. (2023); Stephan et al. (2024); Saha et al. (2025), wherein LLMs evaluate and generate reasoning data. However, this method often introduces bias Guan et al. (2025). Generating labelled training data for step-by-step reasoning remains resource-intensive as well, requiring substantial human effort for annotation, and automatic annotation methods have shown limited success, as noisy reward scores often impede their efficacy Luo et al. (2024); Wang et al. (2024); Chen et al. (2024). Recently, iterative techniques like Monte Carlo Tree Search (MCTS) have been applied to refine reasoning steps and achieve accurate solutions Guan et al. (2025); Park et al. (2024). While MCTS enhances reasoning quality, it is computationally expensive due to extensive rollouts, limiting its scalability for large-scale applications.
@@
Building on this foundation, we introduce Theorem Prover as a Judge (TP-as-a-Judge), a novel framework for generating synthetic data in mathematical reasoning tasks. TP-as-a-Judge is integrated with a theorem prover, leveraging its step-by-step verification to ensure logical correctness. We then propose Reinforcement Learning from Theorem Prover Feedback (RLTPF), a novel approach in which the theorem prover replaces human annotators by providing formalised feedback. By applying TP-as-a-Judge and RLTPF, we achieve competitive performance compared to models fine-tuned on substantially larger datasets in GSM8K and AIME 2024 dataset. This demonstrates the significance of intermediate reasoning in data quality and enhances the data efficiency of model training.
@@ 2 Theorem Prover as a Judge
As shown in Figure 2, TP-as-a-Judge consists of two key stages: LLM Data Generation, where LLMs generate questions and answers; Theorem Prover Verification, where a theorem prover assesses the validity of the model’s reasoning on the generated data and proceeds to RLTPF, leveraging verification results to refine the model through Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO) Rafailov et al. (2024).
@@ 2.2 Theorem Prover Verification
After generating the samples, we use gpt-4o and chatgpt-4o-latest Hurst et al. (2024) as two primary models to generate solutions for the synthetic questions. Subsequently, gpt-4o is used for autoformalisation, after which the outputs are passed to the theorem prover for validation. The Theorem Prover Verification process is divided into two stages: Question and Answer formalisation.
@@ 2.2.1 Question Formalisation
To address the formalisation errors raised by Yang et al. (2024b), which highlight the inherent difficulty of question autoformalisation due to the absence of definitive metrics for assessing correctness, we introduce the third phase , Auto-Informalisation, where the formalised statement is translated back into natural language via GPT-4o. This is followed by the final phase , Alignment Check, which employs a second gpt-4o verification step to ensure alignment between and . This serves as a consistency check, confirming the correctness of the autoformalisation process.
@@ 2.2.3 Iterative Autoformalisation
While the formalisation process may introduce errors in initial iterations, we propose iterative autoformalisation, a novel approach in which the model learns from its errors during the autoformalisation process, shown in Algorithm 1. Our approach concentrates on mathematical autoformalisation, correcting each individual step, prompting the model to revise and reverify the autoformalisation using theorem prover. This iterative process reflects the nature of formalisation, as even human experts require multiple attempts to correctly formalise complex problems. The model may need to experiment with different definitions, refine assumptions, and even attempt partial proofs to identify missing conditions. Our results show that iterative refinement improves the execution rate of the Lean prover formalisation code from 60% to 87% after up to five iterations. This result demonstrates the efficacy of iterative refinement in overcoming challenges associated with theorem prover formalisation.
@@ 2.3 Reinforcement Learning from Theorem Prover Feedback (RLTPF)
In conventional RLHF, human evaluators provide structured feedback to help the model distinguish between valid and invalid reasoning. By contrast, RLTPF replaces human evaluators with a theorem prover, which verifies each formalised reasoning step. In our framework, the theorem prover categorises training data based on verification outcomes, ensuring the model learns from both validated reasoning patterns and cases requiring refinement. Specifically, when both the LLMs generate answers verified by the theorem prover, the corresponding dataset is used for Supervised Fine-Tuning (SFT). When one LLM produces a correct response while the other generates an incorrect one, the paired dataset is utilised for Direct Preference Optimisation (DPO). The model then trained with SFT on verified data before undergoing further optimisation with DPO using the paired dataset. Details for the dataset allocation is shown in Table 1.
@@ 3 Experimental Setup
| Model | MultiArith | SVAMP | GSM8K | GSM-Symbolic | AQUA | MATH-500 | AIME 2024 | OlympiadBench |
@@
To demonstrate the effectiveness and generalisability of our approach, we apply it across a variety of LLMs, including Llama-2-7b-chat-hf Touvron et al. (2023), Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct Dubey et al. (2024), OLMo-2-1124-7B-Instruct OLMo et al. (2024), and Mistral-7B-Instruct-v0.3 Jiang et al. (2023). These models were selected as representative general-purpose single-task language models (SLMs) without specific specialisation in mathematical reasoning tasks. As part of our training pipeline, we employ Low-Rank Adaptation (LoRA; Hu et al. 2021) as our fine-tuning baseline to improve computational efficiency. Details on hyperparameters are provided in Appendix B.
@@
We compare GSM8K and AIME dataset across existing models, including DeepSeekMath Shao et al. (2024), fine-tuned on an extensive dataset comprising 120B mathematical tokens; NuminaMath Li et al. (2024), fine-tuned on over 860k mathematical problem-solution pairs using the DeepSeek-Math-7B-Base model; OpenMath2 Toshniwal et al. (2024), fine-tuned Llama-3.1-8B-Base with a significantly larger corpus of 14M problem-solution pairs; and Mathstral-7B The Mistral AI Team (2024).
@@
We report Pass@1 accuracy as the primary evaluation metric for all baselines. The answers were parsed based on an exact match. For answers where exact matching failed, we parsed the numerical values from the model’s final response and compared them against the ground truth. For the MATH-500, AIME, and OlympiadBench datasets, we adopt a similar evaluation strategy to Leang et al. (2024). We use gpt-4o-mini to assess answer alignment by comparing the last two output sentences to the ground truth, generating a numerical score without revealing the question. Further details can be found in Appendix D.
@@ 4.1 Main Results
Among Llama models (Llama-2-7B, Llama-3.2-3B, and Llama-3.1-8B), we observe that while CoT improves accuracy over the standard baseline, TP-as-a-Judge further enhances performance. In particular, Llama-3.1-8B shows a notable increase in GSM-Symbolic (55.60% 58.20%) and AQUA (53.54% 57.09%). Llama-2-7B exhibits substantial improvements in MultiArith (39.44% 59.44%), SVAMP (38.00% 44.00%), and GSM8K (10.33% 16.70%), while Llama-3.2-3B demonstrates an improvement in GSM-Symbolic (55.80% 59.60%). Our empirical findings suggest that the incremental gains from additional model training diminish as models progress from Llama-2 to Llama-3.1 to Llama-3.2. Nevertheless, these results indicate that synthetic data derived solely from GSM8K and MATH generalises well across a range of benchmarks. Notably, even on AQUA, a multiple-choice task performance is enhanced through the inclusion of synthetic data.
@@
Overall, our results demonstrate the potential of TP-as-a-Judge in ensuring data quality and the importance of intermediate reasoning in mathematical problem-solving. Although SIMQA comprises only 3,508 samples, it outperforms models fine-tuned with larger samples of standard synthetic data, highlighting the significance of verifying intermediate reasoning. Increasing the dataset size could potentially improve model performance.
@@ 4.2 Analysis
We then randomly sampled 54 successfully executed cases for manual annotation and evaluate the performance of GPT-4o, o1, and o3-mini on these annotated questions. As shown in Table 5, while the models achieve high accuracy on the GSM8K dataset, their performance declines significantly when answering synthetic questions. This highlights a significant gap in the models’ ability to handle self-generated questions, demonstrating the critical need of theorem prover for verification, and the potential gap of existing LLMs in mathematical reasoning. Detailed examples of manual annotations can be found in Appendix H.
@@ 4.3 Robustness to Iterative Autoformalisation
Given the strong performance of iterative autoformalisation, we explore another key aspect: How many iterations are required for the model to improve its autoformalisation? To do this, we analyse 356 samples from the dataset. We find that approximately 40% require iterative refinement to enhance formalisation Appendix E. Notably, most successful autoformalisations that require iterative autoformalisation are achieved by the third iteration, indicating that multiple rounds of error correction significantly benefit the model. In a minority of cases, more than three iterations are necessary to reach a correct autoformalisation. Instances requiring more iterations often fail, even after seven attempts. By leveraging verification-based self-correction, the model improves its autoformalisation through an iterative cycle. This demonstrates the importance of structured feedback in enhancing the robustness of formal proofs, ensuring that logical errors are systematically identified and resolved.
@@ 4.4 Effectiveness of RLTPF in TP-as-a-Judge
As shown in Table 6, performing SFT using all possible training instances achieves the highest accuracy among SFT, as it utilises a large number of samples. Meanwhile, using only rejected instances results in lower accuracy compared to using only verified instances. This implies that the Theorem Prover provides an effective inductive bias for learning, but SFT fails to fully leverage it.
@@ 5 Related Work
Advancements in mathematical reasoning for LLMs increasingly rely on high-quality CoT datasets, often distilled from frontier models Guo et al. (2025); Wang et al. (2023); Gou et al. (2023b); Luo et al. (2024); Yu et al. (2023); Tang et al. (2024), such as NuminaMath Li et al. (2024) and OpenMathInstruct Toshniwal et al. (2024). Nevertheless, even solvable problems may contain error-prone intermediate reasoning steps, which are inherently challenging to detect. Although rejection sampling methods Yuan et al. (2023); Brown et al. (2024) can improve data quality by filtering out less reliable outputs, they do not guarantee the correctness of intermediate steps. Consequently, the benefits of scaling CoT datasets exhibit diminishing returns, with performance gains nearing saturation. For instance, OpenMathInstruct reported only a 3.9% improvement on the MATH dataset despite an 8× increase in dataset size. Recently, STaR Zelikman et al. (2022), Lean-STaR Lin et al. (2024) and rStar-Math Guan et al. (2025), have been proposed. These approaches rely on generating multiple rollouts and trajectories for verification and synthesis.
Theorem Proving & Autoformalisation
Modern formal mathematics environments typically centre on theorem provers, such as Lean Jiang et al. (2024); Lin et al. (2024), Isabelle Zhou et al. (2024); Xin et al. (2023) and Coq Huet et al. (1997). These systems have been widely used to verify complex mathematical results, including the Liquid Tensor Experiment Castelvecchi (2021), the formalisation of the PER conjecture Gowers et al. (2023), and efforts to formalise graduate-level number theory Eberl et al. (2024). The Draft-Sketch-Prove Jiang et al. (2022) approach enhances language models’ formal proving abilities by generating informal proofs, translating them into formal sketches, and completing them with symbolic tools.

### v2

@@ Abstract
The demand for synthetic data in mathematical reasoning has increased due to its potential to enhance the mathematical capabilities of large language models (LLMs). However, ensuring the validity of intermediate reasoning steps remains a significant challenge, affecting data quality. While formal verification via theorem provers effectively validates LLM reasoning, the autoformalisation of mathematical proofs remains error-prone. We introduce iterative autoformalisation, an approach that iteratively refines theorem prover formalisation to mitigate errors, thereby increasing the execution rate on the Lean prover from 60% to 87%. Building upon that, we introduce Theorem Prover as a Judge (TP-as-a-Judge), a method that makes use of theorem prover formalisation to rigorously assess LLM intermediate reasoning, effectively integrating autoformalisation with synthetic data generation. Finally, we present Reinforcement Learning from Theorem Prover Feedback (RLTPF), a framework that replaces human annotation with theorem prover feedback in Reinforcement Learning from Human Feedback (RLHF). Across multiple LLMs, applying TP-as-a-Judge and RLTPF improves benchmarks with only 3,508 samples, achieving 5.56% accuracy gain on Mistral-7B for MultiArith, 6.00% on Llama-2-7B for SVAMP, and 3.55% on Llama-3.1-8B for AQUA.111Code is available at https://github.com/joshuaongg21/RLTPF.
@@ 1 Introduction
Recent studies have shown that large language models (LLMs) are increasingly capable of tackling mathematical problems Dubey et al. (2024); Jaech et al. (2024); Yang et al. (2024a); Guo et al. (2025). Techniques such as Chain of Thought (CoT) aim to enhance the reasoning process by breaking complex problems into intermediate steps Wei et al. (2022); Kojima et al. (2022). Although these methods have achieved success on benchmarks like GSM8K Cobbe et al. (2021), they are prone to significant flaws. Specifically, LLM-generated CoT reasoning steps may often include errors, leading to incorrect solutions even when the final answer appears plausible Leang et al. (2024).
@@
Parallel to these strategies, synthetic data generation emerged as a promising approach to enhance mathematical reasoning Xin et al. (2024); Lu et al. (2024b). A commonly used verification method is LLM-as-a-Judge Zheng et al. (2023); Stephan et al. (2024); Saha et al. (2025), wherein LLMs evaluate and generate reasoning data. However, this method often introduces bias Guan et al. (2025). Generating labeled training data for step-by-step reasoning remains resource-intensive as well, requiring substantial human effort for annotation, and automatic annotation methods have shown limited success, as noisy reward scores often impede their efficacy Luo et al. (2024); Wang et al. (2024c); Chen et al. (2024). Recently, iterative techniques like Monte Carlo Tree Search (MCTS) have been applied to refine reasoning steps and achieve accurate solutions Guan et al. (2025); Park et al. (2024). While MCTS enhances reasoning quality, it is computationally expensive due to extensive rollouts, limiting its scalability for large-scale applications.
@@
Building on this foundation, we introduce Theorem Prover as a Judge (TP-as-a-Judge), a novel framework for generating synthetic data in mathematical reasoning tasks. TP-as-a-Judge is integrated with a theorem prover, leveraging its step-by-step verification to ensure logical correctness. We then propose Reinforcement Learning from Theorem Prover Feedback (RLTPF), a novel approach in which the theorem prover replaces human annotators by providing formalised feedback. By applying TP-as-a-Judge and RLTPF, we achieve competitive performance in GSM8K and AIME 2024 datasets compared to models fine-tuned on larger datasets. This demonstrates the significance of intermediate reasoning in data quality and enhances the data efficiency of model training.
@@ 2 Theorem Prover as a Judge
As shown in Figure 2, TP-as-a-Judge consists of two key stages: LLM Data Generation, where LLMs generate questions and answers; Theorem Prover Verification, where a theorem prover assesses the validity of the model’s reasoning on the generated data and proceeds to RLTPF, leveraging verification results to refine the model through Supervised Fine-Tuning (SFT) and Direct Preference Optimization (DPO; Rafailov et al. 2023).
@@ 2.2 Theorem Prover Verification
After generating the samples, we use gpt-4o and chatgpt-4o-latest Hurst et al. (2024) as two primary models to generate solutions for the synthetic questions. Subsequently, gpt-4o is used for autoformalisation, after which the outputs are passed to the theorem prover for validation. The Theorem Prover Verification process is divided into two stages: Question and Answer formalisation. A detailed example of the process is shown in Appendix K.
@@ 2.2.1 Question Formalisation
To address the formalisation errors raised by Yang et al. (2024b), which highlight the inherent difficulty of question autoformalisation due to the absence of definitive metrics for assessing correctness, we introduce the third phase , Auto-Informalisation, where the formalised statement is translated back into natural language via gpt-4o. This is followed by the final phase , Alignment Check, which uses a second gpt-4o verification step to ensure alignment between and . This serves as a consistency check, confirming the correctness of the autoformalisation process.
@@ 2.2.3 Iterative Autoformalisation
We propose iterative autoformalisation, an approach in which the model learns from its errors during the autoformalisation process (Algorithm 1). Our approach concentrates on mathematical autoformalisation, correcting each individual step, prompting the model to revise and reverify the autoformalisation using theorem prover. This iterative process reflects the nature of formalisation, as even human experts require multiple attempts to correctly formalise complex problems. The model may need to experiment with different definitions, refine assumptions, or attempt partial proofs to identify missing conditions. Iterative refinement improves the execution rate of the Lean prover formalisation from 60% to 87% after up to five iterations.
@@ 2.3 Reinforcement Learning from Theorem Prover Feedback (RLTPF)
In conventional RLHF, human evaluators provide structured feedback to help the model distinguish between valid and invalid reasoning. By contrast, RLTPF replaces human evaluators with a theorem prover, which verifies each formalised reasoning step. In our framework, the theorem prover categorises training data based on verification outcomes, ensuring the model learns from both validated reasoning patterns and cases requiring refinement. Specifically, when both the LLMs generate answers verified by the theorem prover, the corresponding dataset is used for Supervised Fine-Tuning (SFT). When one LLM produces a correct response while the other generates an incorrect one, the paired dataset is used for Direct Preference Optimisation (DPO). The model is then trained with SFT on verified data before undergoing further optimisation with DPO using the paired dataset. Details for the dataset allocation are shown in Table 1.
@@ 3 Experimental Setup
| Model ↓ Tasks → | MultiArith | SVAMP | GSM8K | GSM-Symbolic | AQUA | MATH-500 | AIME 2024 | OlympiadBench |
@@
To demonstrate the effectiveness and generalisability of our approach, we apply it across a variety of LLMs, including Llama-2-7b-chat-hf Touvron et al. (2023), Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct Dubey et al. (2024), OLMo-2-1124-7B-Instruct OLMo et al. (2025), and Mistral-7B-Instruct-v0.3 Jiang et al. (2023a). These models were selected as representative general-purpose single-task language models (SLMs) without specific specialisation in mathematical reasoning tasks. As part of our training pipeline, we use Low-Rank Adaptation (LoRA; Hu et al. 2022) as our fine-tuning baseline to improve computational efficiency. Details on hyperparameters are provided in Appendix B.
@@
We compare GSM8K and AIME dataset across existing models, including DeepSeekMath Shao et al. (2024), fine-tuned on an extensive dataset comprising 120B mathematical tokens; NuminaMath Li et al. (2024a), fine-tuned on over 860k mathematical problem-solution pairs using the DeepSeek-Math-7B-Base model; OpenMath2 Toshniwal et al. (2024), fine-tuned Llama-3.1-8B-Base with a significantly larger corpus of 14M problem-solution pairs; and Mathstral-7B The Mistral AI Team (2024).
@@
We report Pass@1 accuracy as the primary evaluation metric for all baselines. The answers were parsed based on an exact match. For answers where exact matching failed, we parsed the numerical values from the model’s final response and compared them against the ground truth. For the MATH-500, AIME, and OlympiadBench datasets, we adopt a similar evaluation strategy to Leang et al. (2024). We use gpt-4o-mini to assess answer alignment by comparing the last two output sentences to the ground truth, generating a numerical score (1 or 0) without revealing the question. Further details can be found in Appendix D.
@@ 4.1 Main Results
Among Llama models (Llama-2-7B, Llama-3.2-3B, and Llama-3.1-8B), we observe that while CoT improves accuracy over the standard baseline, TP-as-a-Judge further enhances performance. In particular, Llama-3.1-8B shows a notable increase in GSM-Symbolic (55.60% 58.20%) and AQUA (53.54% 57.09%). Llama-2-7B exhibits substantial improvements in MultiArith (39.44% 59.44%), SVAMP (38.00% 44.00%), and GSM8K (10.33% 16.70%), while Llama-3.2-3B demonstrates an improvement in GSM-Symbolic (55.80% 59.60%). Our empirical findings suggest that the incremental gains from additional model training diminish as models progress from Llama-2 to Llama-3.1 to Llama-3.2. Nevertheless, these results indicate that synthetic data derived solely from GSM8K and MATH generalises well across a range of benchmarks. Even on AQUA, multiple-choice task performance is enhanced through the inclusion of synthetic data.
@@
Overall, our results demonstrate the potential of TP-as-a-Judge in ensuring data quality and the importance of intermediate reasoning in mathematical problem solving. Although SIMQA comprises only 3,508 samples, it outperforms models fine-tuned with larger samples of standard synthetic data, highlighting the significance of verifying intermediate reasoning. Increasing the dataset size could potentially improve model performance.
@@ 4.2 Analysis
We then randomly sampled 54 successfully executed cases for manual annotation and evaluate the performance of gpt-4o, o1, and o3-mini on these annotated questions. As shown in Table 5, while the models achieve high accuracy on the GSM8K dataset, their performance declines significantly when answering synthetic questions. This highlights a significant gap in the models’ ability to handle self-generated questions, demonstrating the importance of theorem prover for verification, and the potential gap of existing LLMs in mathematical reasoning. Details of autoformalisation breakdown can be referred to Appendix 11, and detailed examples of manual annotations can be found in Appendix I.
@@ 4.3 Robustness to Iterative Autoformalisation
Given the strong performance of iterative autoformalisation, we explore another key aspect: How many iterations are required for the model to improve its autoformalisation? To do this, we analyse 356 samples from the dataset. We find that approximately 40% require iterative refinement to enhance formalisation (Appendix E). In addition, most successful autoformalisations that require iterative autoformalisation are achieved by the third iteration, indicating that multiple rounds of error correction significantly benefit the model. In a minority of cases, more than three iterations are necessary to reach a correct autoformalisation. Instances requiring more iterations often fail, even after seven attempts. By leveraging verification-based self-correction, the model improves its autoformalisation through an iterative cycle. This demonstrates the importance of structured feedback in enhancing the robustness of formal proofs, ensuring that logical errors are systematically identified and resolved.
To further analyse the robustness of iterative autoformalisation, we manually inspected ten randomly sampled examples. We observed that errors were typically caused by an incorrect theorem or tactic generated by Lean, or by the model producing invalid Lean syntax, and generally required one or two further iterations to rectify. Feedback from Lean via iterative autoformalisation allows the model to refine both the syntax and the choice of theorem, thereby correctly verifying each reasoning step. This potentially enables the model to analyse reasoning in terms of intermediate steps rather than as a whole, and to verify the intermediate steps of complex proofs accurately.
| Iteration | Average Sequence Length |
| 1 | 731.2 |
| 2 | 869.1 |
| 3 | 937.6 |
| 4 | 1134.6 |
| 5 | 1343.0 |
Based on Table 6, we can conclude that the number of iterations exhibits a strong positive correlation with sequence length (measured in tokens), since more complex reasoning naturally produces longer token sequences Muennighoff et al. (2025). As the iteration count increases, the method accommodates increasingly intricate reasoning cases, which may mitigate bias towards simpler instances.
@@ 4.4 Effectiveness of RLTPF in TP-as-a-Judge
As shown in Table 7, performing SFT using all possible training instances achieves the highest accuracy among SFT, as it uses a large number of samples. Meanwhile, using only rejected instances results in lower accuracy compared to using only verified instances. This implies that the Theorem Prover provides an effective inductive bias for learning, but SFT fails to fully leverage it.
@@ 5 Related Work
Advancements in mathematical reasoning for LLMs increasingly rely on high-quality CoT datasets, often distilled from frontier models Guo et al. (2025); Wang et al. (2024b); Gou et al. (2024b); Luo et al. (2024); Yu et al. (2024); Tang et al. (2024), such as NuminaMath Li et al. (2024a) and OpenMathInstruct Toshniwal et al. (2024). Nevertheless, even solvable problems may contain error-prone intermediate reasoning steps, which are inherently challenging to detect. Although rejection sampling methods Yuan et al. (2023); Brown et al. (2024) can improve data quality by filtering out less reliable outputs, they do not guarantee the correctness of intermediate steps. Consequently, the benefits of scaling CoT datasets exhibit diminishing returns, with performance gains nearing saturation. For instance, OpenMathInstruct reported only a 3.9% improvement on the MATH dataset despite an 8× increase in dataset size. Recently, STaR Zelikman et al. (2022), Lean-STaR Lin et al. (2024) and rStar-Math Guan et al. (2025), have been proposed. These approaches rely on generating multiple rollouts and trajectories for verification and synthesis.
Theorem Proving and Autoformalisation
Modern formal mathematics environments typically centre on theorem provers, such as Lean Jiang et al. (2024); Lin et al. (2024), Isabelle Zhou et al. (2024); Wang et al. (2024a) and Coq Huet et al. (1997). These systems have been widely used to verify complex mathematical results, including the Liquid Tensor Experiment Castelvecchi (2021), the formalisation of the PER conjecture Gowers et al. (2023), and efforts to formalise graduate-level number theory Eberl et al. (2024). The Draft-Sketch-Prove Jiang et al. (2023b) approach enhances language models’ formal proving abilities by generating informal proofs, translating them into formal sketches, and completing them with symbolic tools.